During one 5-day period, Officer Fernandez drove his patrol motorcycle
225 miles. If he drove 85 miles on one day, how many miles did he
average on each of the other days?

Answer by Theo(13342)

he rode 225 miles in 5 days.
on one day, he drove 85 miles.
that means he drove 225 - 85 = 140 miles on the other 4 days.
140 / 4 = an average of 35 miles on each of the other days.
the total miles he traveled would be 85 + 4 * 35 = 85 + 140 = 225.